OpenStudy (anonymous):

segment AD has a midpoint B, and BD has a midpoint C. If A has coordinates (-5,4) and D is (3,0) what id coordinate c?

OpenStudy (amistre64):

it looks like it would be 3/4 from A to D, maybe usually I add the ends and divide in half to get the average point, the middle point, the mid point. but then we would have to do the same but with the midpoint as the end \[\frac{\frac{a+d}{2}+d}{2}\] \[\frac{\frac{a+d+2d}{2}+d}{2}\] \[\frac{\frac{a+d+2d}{2}}{2}\] \[\frac{a+d+2d}{4}\] \[\frac{a+3d}{4}\] might do it

OpenStudy (amistre64):

A(-5,4) +3D (9, 0) ---------- (4,4)/4 = (1,1) if i thought it out correctly